# A Hybrid ABM-PDE Framework for Real-World Infectious Disease Simulations

### v1

@@ Abstract
The coupling mechanism ensures consistency at the model interface: agents crossing from the ABM into the PDE domain are removed and represented as density contributions at the corresponding grid node, while surplus density in the PDE domain is used to generate agents with plausible trajectories derived from mobile phone data.
@@ 1 Introduction
The complete model implementation and the processed infection data from the Robert Koch Institute (RKI) website are available on Zenodo111https://zenodo.org/records/15100902. Mobility data is available from the authors upon reasonable request.
@@ 2.1.2 Extension with Health States
requiring hospitalization (after being symptomatic before)
@@
requiring hospitalization (after being in critical state before)
@@
[14]. Further, we have the constant transition rates for each rule , characteristic function , and an adjacency matrix
@@ 2.2 Part 2: Brandenburg – ABM Based on Mobile Phone Trajectories
Each agent is either located inside a facility – such as home, workplace, school or leisure facility – or commuting between facilities. While commuting, agents are assumed to be isolated and cannot transmit infections. Within facilities, agents can only interact with others inside the same facility and category. This means, infection transmission can occur only if an infectious and a susceptible agent are present in the same location of the same facility type. Health status updates happen at every time step, except for susceptible agents for which the health status is updated only upon leaving a facility. Transitions in health states follow the same rules as described in Section 2.1.2.
@@ 2.4 Coupling
To facilitate the exchange between ABM and PDE in our hybrid model, the event data from the mobility dataset is sliced into chunks corresponding to individual time steps. Artificial events are introduced during this slicing to track the last known positions of all agents, allowing for consistent updates during each time step to exchange information between the models. For simplicity, we used the same strategy for the full-ABM, which may introduce some computational overhead. However, it offers the practical advantage of reduced memory usage by focusing only on the active time window. For visualization and boundary checking, each agent’s exact position is calculated at the end of every time step.
@@ 2.4.1 Agent Enters PDE Domain
When an agent leaves the ABM domain and enters the PDE domain, the agent is removed from the ABM and their contribution is added to the PDE model. To do this, we project the agent’s (continuous) spatial position onto the closest grid cell of the PDE mesh and increase the corresponding compartment density at that location. The amount added is the density equivalent of a single individual, which we precompute for each grid cell. This procedure ensures consistent mass transfer across the model boundary and preserves the total population in the hybrid model.
We assume that a single susceptible individual is to be added at a single grid point , meaning that the density changes only at that specific location. This update is applied at the end of the time step, after the PDE solution has already been computed for all grid points . We denote by the PDE solution at time before adding the contribution from agent transitions. With these assumptions, the following equation is satisfied, ensuring that the total added density corresponds to one individual:
@@
2.4.2 PDE Person Enters Domain of ABM
To simulate agents transitioning from the PDE domain (Berlin) into the ABM domain (Brandenburg), we precompute the expected number of individuals present in the PDE model for each hour of the day. If the time step corresponds to one hour, this allows us to determine how many individuals should leave the PDE model and be represented as agents in the ABM.
The health status of the transitioning individuals is sampled based on the current distribution of the PDE compartment densities. To update the PDE densities, we first attempt to uniformly reduce the relevant compartments across all grid cells. If this leads to any density value becoming negative – which means that more individuals are removed than are available at certain locations – we switch to an alternative approach: We calculate the number of individuals that can be removed from each grid point, and store the corresponding grid indices. These indices are then shuffled, and a subset is randomly selected to determine the locations from which the transitions occur. The densities at the selected grid points are then reduced accordingly.
When an individual enters the ABM domain, we randomly pick a free agent ID, and the corresponding plan from the mobile phone-based event data is activated. This plan defines the agent’s activities and locations for the current day. Occasionally, the agent’s first or final position from that plan might be inside Berlin, i.e. outside of the ABM domain. These cases are not explicitly handled but occur infrequently in practice.
@@ 3 Parameter Identification
For the ABM, we have the calibration parameter corresponding to the infection rate in a compartmental model [22]. All final transition rates can be found in the next Section 4 in Tables 3 and 4.
@@ 4 Implementation
For fitting the constant part of the infection rate in the PDE model (3), we used a simple grid search. We compared the mean absolute errors of five simulation runs against the target data for different parameter values. For the first time interval (March 2, 2020, to March 15, 2020), we initially tested , while for the second time interval (March 16, 2020, to April 28, 2020), we considered (see Table 1). We tested all combinations together, as the best choice for the first interval may minimize the error in that interval but could be far off for the second interval. The second interval begins with the implementation of school closures and is characterized by a significant change in the trajectory of the infection curve. The smallest errors were achieved with the parameters and (see Table 1). The parameter fitting for the scalar factor of the ABM infection rate was also performed using a simple grid search (see Table 2). All final transition rates can be found in Tables 3 and 4.
@@
March 16 - April 28, 2020 | |||||
29.37 | |||||
March 2, 2020 - April 28, 2020 | |||
41.25 |
@@
We created the modeling domain of Berlin using the boundary grid nodes of OpenDataLab [10]. The triangulation of the modeling domain Berlin was performed using Triangle [26], while the hybrid ABM-PDE systems were solved using the finite element method implemented in the Kaskade7 software [13], with the Dune interface. For the coupling of the hybrid model, we use the nanoflann library [7] for efficient KD-tree construction in our implementation. Specifically, we use it to find the closest grid point to the position of an agent. The visualization of our simulation results was done using the matplotlib package in Python 3.9 and the grid was visualized using ParaView [27].
We computed the 7-day average of Covid-19 data from the RKI [16] and used it as our target data for parameter fitting, evaluation, and for our initial conditions. Our ABM is a streamlined version of EpiSim [1], implemented following[22, 23], with the exclusion of infection dynamics outside of facilities and without the use of weather data.
@@
While prior work indicates that the outcomes of the 25% and 100% population models are largely similar, we explicitly examine both in this paper. Since the calibration parameter was fitted to the 25% model, discrepancies are expected when applying it to the 100% model. In particular, the 100% model is likely to exhibit higher error, as no additional parameter optimization was performed. We measure both the average error and the average runtime to better understand the performance variations between the models. Particularly for the ABM and the hybrid model, we aim to assess how well the rates from the 25% model can be transferred to the 100% model. This will help determine whether parameter optimization can be performed using faster models and, for example, enable predictions for the entire population. Additionally, it will be interesting to compare whether the hybrid model with 100% of the population or the ABM with 25% of the population is faster, and which model potentially results in a lower error.
Implementing the model for 100% of the population results in changes to several data files. The event data file will be updated to include more people with events, the facility data file will expand to list additional facilities, and the maximum number of agents per facility in a given category will increase as more agents visit the same locations. Consequently, more facilities will be included overall. In [22, 23], it was mentioned that needs to be adjusted for the entire population. However, we will not modify this parameter, as other factors, such as room size, will also change, effectively counteracting the need to scale . The calibration parameter will remain unchanged. The infection rate in the ABM depends on multiple factors, with contact intensity being the most affected. This is primarily due to changes in the maximum number of agents per facility within a given category when adjusting the dataset size.
It is worth noting that the infection rate in the ABM depends on multiple factors. Among them, contact intensity is most affected by scaling, since the number of agents per facility changes with dataset size. The infection rate in the PDE model, however, must be explicitly rescaled. This is due to its nonlinear structure: it includes terms that represent pairwise interactions (e.g., between susceptibles and infectious), which scale quadratically with population size. In contrast, the other transition rates represent individual-level processes and remain unchanged.
@@ 5.2 Number of Runs
The relative difference between consecutive values is used as a threshold to determine the necessary number of runs. With a 1% relative difference between consecutive values, the hybrid model requires 29 runs, while the full-ABM requires 92 runs. Beyond 92 runs, the difference between consecutive values remains below 1%, meaning additional runs (e.g., 93) do not significantly impact the result. This results in a 64-run reduction for the hybrid model.
For a 0.5% relative difference, the hybrid model requires 54 runs, whereas the full-ABM needs 100 runs, leading to a 46-run reduction for the hybrid model.
@@
The cumulative mean running times of 10 runs are displayed in Figs. 5(a) and 5(b) for the full-ABM and the hybrid model, respectively. The hybrid model achieves a threshold of 0.05% within the first two runs, whereas the ABM is initially close to 0.12%. For a threshold of 0.02%, the hybrid model achieves this again within two runs, while the ABM requires 10 runs. The hybrid model reaches 0.01% after just 9 runs, whereas the ABM’s relative difference between the last two cumulative runs remains around 0.016%.
@@
The cumulative absolute mean error for the ABM and hybrid model, based on 100 runs with the full population sample (100%), is shown in Figs. 6(a) and 6(b), respectively. With a 1% relative difference between consecutive values, the hybrid model requires 28 runs, whereas the full-ABM only needs 10. Beyond this point, additional runs (e.g., 11) have little effect, as the relative difference remains below 1%, resulting in an 18-run reduction for the full-ABM.
For a 0.5% relative difference, the hybrid model requires 64 runs, while the full-ABM needs only 30, leading to a 34-run reduction for the full-ABM. Notably, when the threshold is halved, the hybrid model requires about twice as many runs, whereas the ABM’s run count triples. Despite requiring fewer runs, the ABM now shows a clear advantage over the hybrid model, as it remains significantly more efficient in this setting.
We observe a significant decrease in the number of runs required for the ABM when increasing the population from a 25% sample to the full population. Assuming that the ABM for the full population takes four times longer to run than for the partial population, the choice of which sample to use depends on the selected threshold in combination with the number of runs and the hypothetical duration of a single run. For a threshold of 1%, using the full population is clearly more efficient, whereas for a threshold of 0.5%, the smaller sample proves to be the better option.
Interestingly, when examining the hybrid model results, the number of runs required for the full population either increases or remains almost the same compared to the partial sample. In contrast, the ABM shows a noticeable decrease in the required number of runs. Regardless of the chosen threshold, this means that the hybrid model with the smaller population sample is, overall, probably the better choice for saving total simulation time to achieve stable results, assuming that the model for the full population will be slower than for the partial sample. For a proper efficiency comparison between the hybrid model and the ABM, the exact running times per single run are required.
Figs. 7(a) and 7(b) present the cumulative mean running times over 10 runs for the full-ABM and hybrid model, respectively. Again, the hybrid model achieves a threshold of 0.05% with only the first two runs. In contrast, the ABM is close to 0.7%. For a 0.02% relative difference, the hybrid model reaches the threshold with the first two runs, while the ABM requires more than 10 runs.
For a 0.01% relative difference, the hybrid model achieves the desired threshold with just 10 runs. Specifically, the relative difference between the first two cumulative runs is 0.0121%. In contrast, the ABM requires more than 10 runs, with the relative change between the last consecutive runs being 0.0431%.
@@
where and denote the (absolute) mean values of the hybrid and full-ABM models, respectively. Computational time was measured with the parameter setting verbosity=0 in the covid.cpp file, which reduces print outputs, prevents the creation of plots, disables the saving of agent health state and location data, and avoids storing density-related information.
@@
The fitted results for the hybrid model and the full-ABM, along with their corresponding target data, are shown in Figs. 9(a) and 9(b), respectively. Agein, these results represent the mean across 100 simulation runs, with the associated standard deviation displayed in the figures.
The error results for the full-ABM and hybrid model are presented in Table 7, while the runtime for both models is shown in Table 8. The fitting was performed using the 25% population dataset, so we expected higher errors when applying it to the full population dataset. However, this is only the case for the ABM, while the hybrid model does not exhibit a higher error. Hence, the increased error compared to the previous model went from -11.16% to -51.25%.
The runtime of the hybrid model has almost quadrupled, while the runtime of the ABM has increased by more than a factor of four. As a result, the decreased duration compared to the previous model rises from 30.62% to 37.06%.
@@
Now, knowing the least number of runs needed for both thresholds (1% and 0.5%), as well as the runtime for a single run, we can conclude that the total simulation time of the hybrid model for the partial sample is the smallest, requiring at least hours for the 1% threshold and hours for the 0.5% threshold. In the case of the ABM, it depends on the chosen threshold. For the 1% threshold, the full population sample should be used, which requires a total simulation time of at least hours. For the 0.5% threshold, the partial population sample should be used, which requires a total simulation time of at least hours.
@@ 5.4 Extreme Cases
To test whether the coupling works in both directions and to evaluate the model’s stability, we will conduct experiments for extreme cases. People continue to follow their daily activities, and restrictions are imposed as before. The only change we make is to the initial health state of all individuals. All agents are initially exposed and therefore soon become infectious, while the PDE population is initially susceptible, and vice versa. In these experiments, we can observe the influence of commuting individuals from Berlin or Brandenburg on each other’s infection dynamics.
@@
The total number of symptomatic individuals for the ABM contribution and the PDE contribution (3) of the hybrid model are shown in Figs. 10(a) and 11(a), while snapshots of the corresponding experiments can be found next to these plots in Figs. 10(b) and 11(b). For better comparability, we set the maximum of the y-axis to the same value in the total number plots.
Examining the total number of plots, we observe a significant difference in how the states influence each other. In the first experiment (see Fig. 10(a)), the peak number of symptomatic individuals in Berlin (blue) is significantly higher than in Brandenburg, to the extent that the small wave in Brandenburg is barely visible. In the second experiment (see Fig. 11(a)), the peaks have shifted toward each other: the peak in Berlin is lower than in the first experiment, while the peak in Brandenburg is now clearly visible. However, the peak in Brandenburg remains significantly lower than the previous peak in Berlin from the first experiment. It is evident that there is a notable difference depending on which model the entire population is set to exposed or susceptible.
Additionally, we observe the formation of a second wave in Brandenburg in the second experiment, with a sharp decline followed by a significant increase in symptomatic cases between the waves. However, upon closer inspection of the data (see Fig. 10(a)), this pattern is also present in the first experiment in the ABM, where a small second wave, if we may call it that, forms in a similar manner. The increases and decreases in the first experiment in the ABM, where we observe very few cases, seem to be partially driven by the activity data (see Fig. 2). However, this is less the case for the other scenarios, where the waves are more pronounced, and the symptomatic cases are significant. Nevertheless, it is possible that the activity data triggered the formation of the second wave. Interestingly, in the first experiment for the ABM, the number of symptomatic cases even reaches 0.0 between the waves.
The peak in Berlin occurs on day 18, shortly after the second time interval with the lower infection rate begins. It is worth noting that the activity change rates, together with the increasing cases in Berlin, may have contributed to the formation of a second wave. After adjusting the implementation to prevent individuals from jumping out of the PDE domain into the ABM (see Fig. 12), a second wave is still visible, and it appears almost identical in the case of the ABM compared to the previous experiment with full coupling. Hence, the influence of Berlin is small or not clearly visible in this case. Additionally, as the individuals in the PDE domain are required to stay in Berlin, symptomatic cases have slightly increased in Berlin.
@@
Similarly, the total number of symptomatic individuals for the ABM and PDE contributions (3) of the hybrid model are presented in Figs. 13(a) and 14(a). Corresponding experiment snapshots are displayed next to these plots in Figs. 13(b) and 14(b). Again, to ensure comparability, the y-axis maximum has been set to the same value across the total number plots.
The upper plot in Fig. 13(a) appears identical to the upper plot in Fig. 10(a), whereas the lower plot closely resembles the lower plot in Fig. 10(a). The peak of the ABM curve is higher for the partial population, but it is important to note that this result is based on a single run. Once again, the increases and decreases, particularly in the first experiment within the ABM, appear to be partially driven by the activity data (see Fig. 2).
Looking at the total numbers plot of the second experiment (see Fig. 14(a)), it appears similar to the plot for the 25% population sample (see Fig. 11(a)). Indeed, both plots are quite similar, with a slightly higher overall number of symptomatic individuals. Notably, both peaks are somewhat higher but occur on the same days.

### v2

@@ Abstract
The coupling mechanism ensures consistency at the model interface: agents crossing from the ABM into the PDE domain are removed and represented as density contributions, while surplus density in the PDE domain is used to generate agents with plausible trajectories derived from mobile phone data.
@@ 1 Introduction
The complete model implementation and the processed infection data from the Robert Koch Institute (RKI) website are available on Zenodo111https://zenodo.org/records/15976953. Mobility data is available from the authors upon reasonable request.
@@ 2.1.2 Extension with Health States
requiring hospitalization (after being symptomatic)
@@
requiring hospitalization (after being in critical state)
@@
[13]. Further, we have the constant transition rates for each rule , characteristic function , and an adjacency matrix with
@@ 2.2 Part 2: Brandenburg – ABM Based on Mobile Phone Trajectories
Each agent is either located inside a facility – such as home, workplace, school or leisure facility – or commuting between facilities. While commuting, agents are assumed to be isolated and cannot transmit infections. Within facilities, agents can only interact with others inside the same facility and category. This means, infection transmission can occur only if an infectious ( or ) and a susceptible () agent are present in the same location of the same facility type. Health status updates happen at every time step, except for susceptible agents for which the health status is updated only upon leaving a facility. Transitions in health states follow the same rules as described in Section 2.1.2.
@@ 2.4 Coupling
To facilitate the exchange between ABM and PDE in our hybrid model, the event data from the mobility dataset is sliced into chunks corresponding to individual time steps. Artificial events are introduced during this slicing to track the last known positions of all agents, allowing for consistent updates during each time step to exchange information between the models. For simplicity, we used the same strategy for the full-ABM, which may introduce some computational overhead. However, it offers the practical advantage of reduced memory usage by focusing only on the active time window. For visualization and boundary checking in the hybrid model, each agent’s exact position is calculated at the end of every time step.
@@ 2.4.1 Agent Enters PDE Domain
When an agent leaves the ABM domain and enters the PDE domain, the agent is removed from the ABM and their contribution is added to the PDE model. The agent’s health state is read and used to place them into the corresponding compartment of the PDE system. This procedure ensures consistent mass transfer across the model boundary and preserves the total population in the hybrid model.
2.4.2 PDE Person Enters Domain of ABM
To simulate agents transitioning from the PDE domain (Berlin) into the ABM domain (Brandenburg), we precompute the expected number of individuals present in the PDE model for each hour of the day. If the time step corresponds to one hour, this allows us to determine how many individuals should leave the PDE model and be represented as agents in the ABM. For smaller time steps, we interpolate these values accordingly.
The health status of the transitioning individuals is not sampled stochastically but determined deterministically based on the current distribution of the PDE compartment densities. Instead of drawing random health states, we compute the ratio of individuals to be removed from each compartment, proportional to their current densities. These initial ratios are adjusted so that we end up rounding down to whole individuals. Any remaining individuals are assigned to the susceptible compartment, which dominates the distribution. These ratios are then used to proportionally reduce the corresponding densities at each grid point, ensuring that more individuals are removed where the density is higher.
To obtain the total number of individuals, we read the density values at each node and convert the density amount into the corresponding number of people. For this, we compute how much density is equivalent to a single person. We assume that a single susceptible individual is to be added at a single grid point , meaning that the density changes only at that specific location. This update is applied at the end of the time step, after the PDE solution has already been computed for all grid points . We denote by the PDE solution at time before adding the contribution from agent transitions. With these assumptions, the following equation is satisfied, ensuring that the total added density corresponds to one individual:
@@ 2.4.2 PDE Person Enters Domain of ABM
When an individual enters the ABM domain, we randomly pick a free agent ID, and the corresponding plan from the mobile phone-based event data is activated. This plan defines the agent’s activities and locations for the current day. Occasionally, the agent’s first or final position from that plan might be inside Berlin, i.e. outside of the ABM domain. These cases are not explicitly handled but occur infrequently in practice compared to the total population size. As agents enter Berlin during the timestep, they can still interact with other agents and contribute to the infection dynamics if a less appropriate plan was chosen, causing them to be placed in an unintended state.
@@ 3 Parameter Identification
For the ABM, we have the calibration parameter corresponding to the infection rate in a compartmental model [21]. All final transition rates can be found in the next Section 4 in Tables 4 and 5.
@@ 4 Implementation
For fitting the constant part of the infection rate in the PDE model (3), we used a simple grid search. We compared the mean absolute errors of 20 simulation runs against the target data for different parameter values. For the first time interval (March 2, 2020, to March 15, 2020), we initially tested , while for the second time interval (March 16, 2020, to April 28, 2020), we considered (see Table 1). We tested all combinations together, as the best choice for the first interval may minimize the error in that interval but could be far off for the second interval. The second interval begins with the implementation of school closures and is characterized by a significant change in the trajectory of the infection curve. The smallest average error was achieved with the parameters and (see Table 1). The parameter fitting for the scalar factor of the ABM infection rate was also performed using a simple grid search (see Tables 2 and 3). All final transition rates can be found in Tables 4 and 5.
@@
March 16 -
April 28, 2020
| 25.12 | |||||
| March 2, 2020 - April 28, 2020 | |||||
| 27.91 | |||||
| March 2, 2020 - April 28, 2020 | |||||
| 63.77 | |||||
@@
We created the modeling domain of Berlin using the boundary grid nodes of OpenDataLab [9]. The triangulation of the modeling domain Berlin was performed using Triangle [25], while the hybrid ABM-PDE systems were solved using the finite element method implemented in the Kaskade7 software [12], with the Dune interface. The visualization of our simulation results was done using the matplotlib package in Python 3.9 and the grid was visualized using ParaView [26].
We computed the 7-day average of Covid-19 data from the RKI [16] and used it as target data for parameter fitting, evaluation, and to define the initial conditions of the PDE system. For the ABM, the data were scaled by a factor of 20. Due to stochasticity in the ABM, infections can die out more easily in the early phase; therefore, higher initial values were chosen compared to the official RKI data.
Our ABM is a streamlined version of EpiSim [1], implemented following[21, 22], with the exclusion of infection dynamics outside of facilities and without the use of weather data.
@@
While prior work indicates that the outcomes of the 25% and 100% population models are largely similar, we explicitly examine both in this paper. We measure both the average error and the average runtime to better understand the performance variations between the models. Particularly for the PDE contribution of the hybrid model, we aim to assess how well the rates from the 25% model can be transferred to the 100% model. This will help determine whether parameter optimization can be performed using faster models and, for example, enable predictions for the entire population. Additionally, it will be interesting to compare whether the hybrid model with 100% of the population or the ABM with 25% of the population is faster, and which model potentially results in a lower error.
Implementing the model for 100% of the population results in changes to several data files. The event data file will be updated to include more people with events, the facility data file will expand to list additional facilities, and the maximum number of agents per facility in a given category will increase as more agents visit the same locations. Consequently, more facilities will be included overall. In [21, 22], it was mentioned that needs to be adjusted for the entire population. However, we will not modify this parameter, as other factors, such as room size, will also change, effectively counteracting the need to scale . The calibration parameter will be optimized.
It is worth noting that the infection rate in the ABM depends on multiple factors. Among them, contact intensity is probably the most affected by scaling, since the number of agents per facility changes with dataset size. The infection rate in the PDE model, however, must be explicitly rescaled. This is due to its nonlinear structure: it includes terms that represent pairwise interactions (e.g., between susceptibles and infectious), which scale quadratically with population size. In contrast, the other transition rates represent individual-level processes and remain unchanged.
@@ 5.2 Number of Runs
The relative difference between consecutive values is used as a threshold to determine the necessary number of runs. With a 2% relative difference between consecutive values, the hybrid model requires 13 runs, while the full-ABM requires 77 runs. Beyond 77 runs, the difference between consecutive values remains below 2%, meaning additional runs (e.g., 78) do not significantly impact the result. This results in a 64-run reduction for the hybrid model.
For a 1% relative difference, the hybrid model requires 31 runs, whereas the full-ABM needs 92 runs, leading to a 61-run reduction for the hybrid model.
@@
The cumulative mean running times of 10 runs are displayed in Figs. 5(a) and 5(b) for the full-ABM and the hybrid model, respectively. The hybrid model achieves a threshold of 0.56% within the first two runs, whereas the ABM is initially close to 1.56%. For a threshold of 0.2%, the hybrid model achieves this again within 4 runs, while the ABM requires 5 runs. The hybrid model reaches 0.01% after just 9 runs, whereas the ABM’s relative difference between the last two cumulative runs remains around 0.16%.
@@
The cumulative absolute mean error for the ABM and hybrid model, based on 100 runs with the full population sample (100%), is shown in Figs. 6(a) and 6(b), respectively. With a 2% relative difference between consecutive values, the hybrid model requires only 4 runs, whereas the full-ABM needs 16. Beyond this point, additional runs (e.g., 17) have little effect, as the relative difference remains below 2%, resulting in a 12-run reduction for the hybrid model.
For a 1% relative difference, the hybrid model requires only 33 runs, while the full-ABM needs 44, leading to a 11-run reduction for the hybrid model. Notably, when the threshold is halved, the hybrid model requires about eight times as many runs, whereas the ABM’s run count almost triples. Again, the hybrid model shows an advantage over the ABM in terms of the required number of runs for these chosen thresholds. For smaller thresholds, however, this may not always be the case.
We observe a significant decrease in the number of runs required for the ABM when increasing the population from a 25% sample to the full population. Assuming that the ABM for the full population takes four times longer to run than for the partial population, the choice of which sample to use depends on the selected threshold in combination with the number of runs and the hypothetical duration of a single run. For a threshold of 2%, using the full population is clearly more efficient, whereas for a threshold of 1%, the smaller sample proves to be the better option.
Interestingly, when examining the hybrid model results, the number of runs required for the full population either decreases or remains almost the same compared to the partial sample. In contrast, the ABM shows a noticeable decrease in the required number of runs. Regardless of the chosen threshold, this means that the hybrid model with the smaller population sample is, overall, probably the better choice for saving total simulation time to achieve stable results, assuming that the model for the full population will be slower than for the partial sample. For a proper efficiency comparison between the hybrid model and the ABM, the exact running times per single run are required.
Figs. 7(a) and 7(b) present the cumulative mean running times over 10 runs for the full-ABM and hybrid model, respectively. The hybrid model achieves a threshold of 0.08% with only the first two runs. In contrast, the ABM is close to 3.09%. For a 0.02% relative difference, the hybrid model reaches the threshold with the first 6 runs, while the ABM requires more than 10 runs.
For a 0.01% relative difference, the hybrid model achieves the desired threshold with just 10 runs. Specifically, the relative difference between the last two cumulative runs is 0.004%. In contrast, the ABM requires more than 10 runs, with the relative change between the last consecutive runs being 0.21%.
@@
where and denote the (absolute) mean values of the hybrid and full-ABM models, respectively. Computational time was measured with the parameter setting verbosity=0 in the covid.cpp file, which reduces print outputs, prevents the creation of plots, disables the saving of agent health state and location data, and avoids storing density-related information. The parameter setting verbosity=1 prints and stores all information except agent details.
@@
The fitted results for the hybrid model and the full-ABM, along with their corresponding target data, are shown in Figs. 9(a) and 9(b), respectively. Again, these results represent the mean across 100 simulation runs, with the associated standard deviation displayed in the figures.
The error results for the full-ABM and hybrid model are presented in Table 8, while the runtime for both models is shown in Table 9. The fitting was performed using the 25% population dataset for the PDE contribution of the hybrid model, so we already expected higher errors when applying it to the full population dataset. This is indeed the case: the error almost doubled. However, we did not expect the total population dataset, which was also fitted to the full-ABM, to have an error almost double compared to the model fitted to the partial dataset (25%). The increased error compared to the previous model went from -27.25% to -27.44%.
The runtime of the hybrid model and the ABM have increased by almost a factor of five. As a result, the decreased duration compared to the previous model rises from 21.85% to 23.49%.
@@
Now, knowing the least number of runs needed for both thresholds (2% and 1%), as well as the runtime for a single run, we can conclude that the total simulation time is smallest for the partial sample at the threshold 1% and for the full population sample at the threshold 2%. The hybrid model requires at least hours of simulation time at the 2% threshold (full population sample) and hours at the 1% threshold (partial sample). In comparison, the full-ABM requires hours at the 2% threshold (full population sample) and hours at the 1% threshold (partial sample).
@@ 5.4 Extreme Cases
To test whether the coupling works in both directions and to evaluate the model’s stability, we will conduct experiments for extreme cases. People continue to follow their daily activities, and restrictions are imposed as before. The only change we make is to the initial health state of all individuals. In the first experiment, all agents are initially susceptible, while the PDE population is initially exposed except for a single susceptible individual, and the exposed quickly become infectious. In the second experiment, all agents start as exposed, whereas the PDE population is initially susceptible except for one exposed individual. This single individual ensures that the PDE population is initially distributed according to the landscape, and that subsequently entering individuals are added (or removed) based on the current distribution. Without this individual, new entrants would be uniformly distributed across the domain, which would not resolve quickly enough and could destabilize the model. An alternative would be to initialize the distribution of entering individuals directly based on the landscape. In these experiments, we can observe the influence of commuting individuals from Berlin or Brandenburg on each other’s infection dynamics.
@@
The total number of symptomatic individuals for the ABM contribution and the PDE contribution (3) of the hybrid model are shown in Figs. 10(a) and 11(a), while snapshots of the corresponding experiments can be found next to these plots in Figs. 10(b) and 11(b). For better comparability, we set the maximum of the y-axis to the same value in the total number plots, except for the close-up.
Examining the total number of symptomatic individuals in these plots, we observe a significant difference in how the states influence each other. In the first experiment (see Fig. 10(a)), the peak number of symptomatic individuals in Berlin (blue) is significantly higher than in Brandenburg, to the extent that the small wave in Brandenburg is barely comparable. In the second experiment (see Fig. 11(a)), the peaks have shifted toward each other: the peak in Berlin is lower than in the first experiment, while the peak in Brandenburg is now clearly higher. However, the peak in Brandenburg remains significantly lower than the previous peak in Berlin from the first experiment. It is evident that there is a notable difference depending on which model the entire population except one individual is set to exposed or susceptible.
Additionally, we observe the formation of a second wave in Brandenburg in the second experiment, with a sharp decline followed by a significant increase in symptomatic cases between the waves. However, upon closer inspection of the data (see Fig. 10(a)), this pattern is also present in the first experiment in the ABM, where a small second wave, if we may call it that, forms in a similar manner. The increases and decreases in the first experiment in the ABM, where we observe very few cases, seem to be partially driven by the activity data (see Fig. 2). However, this is less the case for the other scenarios, where the waves are more pronounced, and the number of symptomatic cases is significantly higher, or the first wave emerges later. Nevertheless, it is possible that the activity data triggered the formation of the second wave. Interestingly, in the first experiment for the ABM, the number of symptomatic cases even reaches almost 0.0 between the waves.
The peak in Berlin occurs on day 12 (see Fig. 11(a)), shortly before the second time interval with the lower infection rate begins. It is worth noting that the activity change rates, together with the increasing cases in Berlin, may have contributed to the formation of a second wave. After adjusting the implementation to prevent individuals from jumping out of the PDE domain into the ABM (see Fig. 12), a second wave is still visible, and it appears almost identical in the case of the ABM compared to the previous experiment with full coupling. Hence, the influence of Berlin is small or not clearly visible in this case. Additionally, as the individuals in the PDE domain are required to stay in Berlin, symptomatic cases have increased in Berlin.
@@
Similarly, the total number of symptomatic individuals for the ABM and PDE contributions (3) of the hybrid model are presented in Figs. 13(a) and 14(a). Corresponding experiment snapshots are displayed next to these plots in Figs. 13(b) and 14(b). Again, to ensure comparability, the y-axis maximum has been set to the same value across the total number plots, except for the close-up.
The upper and lower plots in Fig. 13(a) appear identical to the upper and lower plots in Fig. 10(a), respectively. The peak of the ABM curve is higher for the full population, but it is important to note that this result is based on a single run. Once again, the increases and decreases, particularly in the first experiment within the ABM, appear to be partially driven by the activity data (see Fig. 2).
Looking at the total numbers plot of the second experiment (see Fig. 14(a)), it appears similar to the plot for the 25% population sample (see Fig. 11(a)). Indeed, both plots are quite similar, with a slightly higher overall number of symptomatic individuals. Notably, both peaks are somewhat higher but occur on the same days (5 and 19).